OpenStudy (anonymous):

A circle has a diameter of 6, an area of b square units, and a circumference of c units what is the value of b+ c The answer the book has is 15 pie

OpenStudy (anonymous):

\[r=d/2=6/2=3\] \[areaofacircle=\pi*r^2\] \[b=\pi*(3)^2\] \[b=9\pi\] \[circumferenceofacircle=2*\pi*r\] \[c=2*\pi*3\] \[c=6\pi\] \[b+c=9\pi+6\pi=15\pi\]

OpenStudy (anonymous):

Ok got it thx